Brandon is 3 times as old as Caitlan. 18 years ago, Brandon was 6 times as old as Caitlan was 18 years ago. How old are they now?
Answer by scott8148(6628)

"Brandon is 3 times as old as Caitlan" __ B=3C

"18 years ago, Brandon was 6 times as old as Caitlan was 18 years ago" __ B-18=6(C-18)

substituting __ 3C-18=6C-108 __ 90=3C